# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v1

@@ V-A Challenges
AiOIR models encounter several challenges that limit their effectiveness in real-world applications. Task conflicts arise from differing objectives in denoising, deblurring, and dehazing, complicating simultaneous optimization and leading to inconsistent performance. Additionally, these models struggle with Out of Distribution (OOD) degradations, as real-world images often exhibit a mix of degradation types that don’t align with training data. The computational demands of current models hinder deployment on resource-constrained devices, necessitating a balance between restoration quality and efficiency. Furthermore, the reliance on large-scale, high-quality labeled datasets poses challenges due to the resource-intensive nature of data acquisition, resulting in generalization issues. Lastly, most models focus on RGB images, while handling high-dimensional data introduces further complexities. Addressing these challenges is essential for improving AiOIR models’ practicality and performance.
@@
Task Conflicts. In AiOIR, task conflicts arise due to differing objectives across various tasks like denoising, deblurring, and dehazing. These conflicts occur because tasks may require opposite optimizations—denoising reduces high-frequency noise, while deblurring enhances high-frequency details. Additionally, data characteristics for different tasks vary, leading to inconsistent performance when training on multiple tasks. Model capacity limitations also hinder the network’s ability to simultaneously handle diverse tasks effectively. To address these challenges, strategies such as multi-branch architectures, balanced multi-task loss functions, phased restoration, and adaptive learning can help mitigate conflicts and improve overall performance in all-in-one restoration models.
@@
Handling OOD Degradations. AiOIR models face significant difficulty in handling highly diverse and unforeseen image degradations, which can be treated as OOD degradations. In real-world scenarios, images may suffer from a combination of different degradations, such as blur, noise, low resolution, and compression artifacts. At the same time, the degree of each degradation type is also varied, and it is likely to be inconsistent with the sample distribution at training time while testing.
@@
Model Complexity and Efficiency. Despite recent advancements in AiOIR, these models are often computationally expensive and complex. Their large size and high computational demands make them impractical for deployment on resource-constrained devices such as mobile phones or embedded systems. Striking a balance between performance and efficiency remains a significant issue, requiring models to maintain high restoration quality without becoming too cumbersome.
@@
Limited High-Quality Data. Many AiOIR models depend on large-scale, high-quality labeled datasets for supervised training, but acquiring these datasets is resource-intensive. Real-world degraded data is often scarce, and the unpredictability of degradations makes it difficult for models to perform well in practical applications. Real-world degradations are typically more complex than those represented in synthetic training datasets, leading to generalization issues. Additionally, domain shifts can hinder performance when models are applied to different image types. To address these challenges, solutions like data augmentation, domain adaptation, self-supervised learning, and transfer learning are being explored, but limitations still exist.
@@
High Dimensional Data. Current IR models primarily focus on 2D images, but handling 3D data and video sequences presents additional challenges. For video restoration, not only does each frame need high-quality restoration, but temporal consistency across frames must also be preserved. This adds complexity and demands more sophisticated methods that can integrate spatial and temporal information simultaneously.
@@ V-B Future Research Directions
In future research, the development of AiOIR methods will focus on several key directions. First, establishing a robust multi-task learning theory will be essential to effectively address task conflicts and optimize information-sharing mechanisms, thereby enhancing system performance. Second, reducing reliance on large-scale labeled datasets by exploring semi-supervised and unsupervised learning methods will improve the model’s adaptability in data-scarce situations. Additionally, designing efficient models suitable for edge computing will make AiOIR more feasible for practical applications. Furthermore, investigating more complex real-world degradation scenarios will drive improvements in model performance across various environments. Finally, integrating large multimodal pretrained models and generative priors will enhance the restoration capabilities by leveraging rich multimodal data. These research directions will lay the foundation for the practicality and flexibility of AiOIR models, enabling them to tackle a broader range of real-world challenges.
Incorporating Robust Multi-Task Learning Theory. The development of a robust multi-task learning theory for AiOIR is still in its early stages, offering significant research opportunities [66, 217]. Key challenges include modeling task conflicts, dynamically assigning task weights, and optimizing the information-sharing mechanisms between tasks to maximize performance without interference [218, 219]. Additionally, understanding and establishing the optimal task sequence or phased restoration process can help improve results, as well as incorporating adaptive task prioritization based on degradation severity. Moreover, a multi-task loss function design that balances conflicting objectives is needed. Advancing MTL theory in these areas will be crucial for creating efficient and generalized AiOIR systems capable of handling a wide variety of real-world degradation scenarios.
@@
Semi-Supervised and Unsupervised Learning Methods. Reducing the dependency on large-scale labeled datasets is essential for the scalability and applicability of AiOIR models. Future research should focus on developing semi-supervised and unsupervised learning approaches that can learn effective representations from unlabeled or partially labeled data [220, 221]. Techniques such as self-supervised learning, contrastive learning, and unsupervised domain adaptation can be leveraged to improve model performance in scenarios where labeled data is scarce or unavailable. By advancing these methods, AiOIR models can become more adaptable to diverse and unforeseen degradations encountered in real-world conditions.
Platform-Aware Model Design and Efficient Approaches. A critical direction is the design of edge models that enhance the applicability of AiOIR in real-world scenarios. This involves creating models that are not only accurate but also efficient in terms of computational resources, making them suitable for deployment on various platforms, including mobile devices and embedded systems [222, 223]. Techniques such as model compression, pruning, quantization, and efficient neural architecture search can be employed to develop lightweight models without significant loss in performance. By focusing on efficient all-in-one approaches, researchers can ensure that AiOIR models are practical for everyday use, bridging the gap between research and real-world application.
@@
Addressing More Practical and Complex Degradations. There is a need to focus on more practical tasks and datasets that reflect the complexities of real-world image degradations [224, 225]. Future research should explore composite and complex degradation scenarios, such as image restoration in nighttime conditions, deraining and dehazing in the dark, and images affected by multiple overlapping degradations rather than isolated, mixed single degradation tasks. Developing and utilizing datasets that capture these challenging conditions will enable models to learn from and be tested on data that closely resembles real-world challenges. This focus will drive the development of AiOIR models that are more robust and effective in practical applications.
@@
Incorporating Large Multimodal Pretrained Models and Generative Priors. Another promising avenue is exploiting large multimodal pretrained models, particularly those incorporating generative models, to enhance AiOIR tasks [226, 227, 228, 229]. Models like CLIP [112] and recent advancements in generative models (e.g., Stable Diffusion [153]) have demonstrated remarkable abilities to capture complex data distributions across multiple modalities. By mining the rich representations and priors from these universal models, AiOIR can benefit from enhanced understanding of image content and context, leading to better restoration in low-level tasks. Integrating these models can help in handling a broader range of degradations and in generating more realistic and high-quality restored images, especially in cases where the degradation is severe or complex.
@@
Leveraging Multi-Modal Information. Most current AiOIR models primarily rely on single-modal image information, such as RGB images, limiting their effectiveness in addressing complex restoration tasks. Future research can focus on integrating multi-modal information—like depth maps, optical flow, and infrared images—into AiOIR frameworks [230, 231, 232]. This integration would provide models with rich contextual and structural insights, enhancing their ability to accurately recover images with varied degradations [233]. For example, infrared data can reveal details hidden in RGB images, enabling more comprehensive restoration. By leveraging multi-modal data, all-in-one models can achieve higher robustness and versatility across various applications, from medical imaging to low-light photography.
Establishment of Standardized Evaluation Protocols and Benchmarks. In contrast to single image restoration tasks, such as super-resolution, image deraining, and image dehaze, where there are well-established benchmark training and testing datasets, the AiOIR task lacks a standardized dataset. To facilitate fair comparison and assessment of AiOIR models, it is important to establish standardized evaluation protocols and comprehensive benchmarks. Creating diverse and representative benchmarks that encompass a wide range of real-world scenarios, including high-resolution image restoration, medical image enhancement, old photo restoration, and adverse weather conditions (e.g., sandstorms, nighttime fog), will allow for a more thorough evaluation of model performance. Standardized benchmarks will help identify the strengths and limitations of different models, fostering progress and encouraging the development of more generalizable AiOIR methods.
@@
Extension to Other Data. In addition to RGB images, extending AiOIR methods to other data, such as video [234, 235], 3D data [235, 236], dynamic event data [237, 238], and hyperspectral data [239], presents significant opportunities for future research. Video restoration requires not only enhancing individual frames but also preserving temporal consistency across frames. Similarly, 3D data restoration involves handling spatial information in multiple dimensions and dealing with depth information. Developing AiOIR models for hyperspectral data requires handling the high dimensionality of this data while maintaining spectral consistency across bands. Developing techniques that effectively integrate spatial, temporal, spectral, dynamic, and 3D information will be vital for applications such as video enhancement, 3D rendering, spectral analysis, and augmented reality . Addressing these challenges will expand the capabilities of AiOIR models, making them more versatile and applicable to a broader range of tasks.
@@ VI Conclusion
In this paper, we present a comprehensive review of recent advancements in all-in-one image restoration (AiOIR), a rapidly emerging domain that integrates multiple types of image degradations into a single framework. Through an in-depth exploration of state-of-the-art models, we emphasize their robust capabilities, diverse architectures, and the rigorous experimentations. By contrasting these models with traditional single-task approaches, we underscore the limitations of the latter in addressing real-world complexities, while highlighting the significant gains in efficiency, adaptability, and scalability offered by AiOIR models.
Our comprehensive taxonomy of existing works offers a multi-dimensional perspective, covering structural innovations, key methodologies such as prompt learning, Mixture-of-Experts (MoE), and the incorporation of multimodal models. We further present an analysis of critical datasets, providing researchers and practitioners a holistic toolkit to better assess the current landscape of AiOIR. Despite the considerable strides made in recent years, challenges persist. Current models still struggle with handling complex and compounded degradations, lack computational efficiency, and do not generalize well in real-world scenarios. We believe that future research will focus on several key areas: the development of more lightweight and efficient architectures, advancements in semi-supervised learning, and expanding the scope of AiOIR models to accommodate multimodal inputs and video data. Additionally, as the field progresses, innovations in cross-modal learning, real-time processing, and interpretability will likely become central to pushing the boundaries of what AiOIR systems can achieve.
In conclusion, AiOIR represents a promising, unified approach to tackling diverse degradation challenges in a more holistic manner. As the field continues to evolve, it holds immense potential for broader applications in real-world contexts, from media enhancement to autonomous systems. We hope that this review not only maps the current state-of-the-art but also inspires further innovations and breakthroughs in the pursuit of more sophisticated, efficient, and versatile AiOIR models.

### v2

@@ V-A Challenges
AiOIR models encounter several challenges that limit their effectiveness in real-world applications. Task conflicts arise from differing objectives in denoising, deblurring, and dehazing, complicating simultaneous optimization and leading to inconsistent performance. Additionally, these models struggle with out-of-distribution (OOD) degradations, as real-world images often exhibit a mix of degradation types that do not align with training data. The computational demands of current models hinder deployment on resource-constrained devices, necessitating a balance between restoration quality and efficiency. Furthermore, the reliance on large-scale, high-quality labeled real-world datasets poses challenges due to the resource-intensive nature of data acquisition, resulting in generalization issues. Lastly, most models focus on RGB images, while handling high-dimensional data introduces further complexities, such as increased dimensionality and temporal consistency requirements. Addressing these challenges is essential for improving AiOIR models’ practicality and performance.
V-A1 Task Conflicts
@@
In AiOIR, task conflicts arise due to differing objectives across various tasks like denoising, deblurring, and dehazing. Tasks may require opposite optimizations—denoising reduces high-frequency noise, while deblurring enhances high-frequency details. Additionally, data characteristics for different tasks vary, leading to inconsistent performance when training on multiple tasks.
V-A2 OOD Degradations
@@
AiOIR models face significant difficulty in handling highly diverse and unforeseen image degradations, which can be treated as OOD degradations. In real-world scenarios, images may suffer from a combination of different degradations, such as blur, noise, low resolution, and compression artifacts. At the same time, the degree of each degradation type varies, and may be inconsistent with the sample distribution at training time while testing.
V-A3 Model Complexity and Efficienc.
@@
Despite recent advancements in AiOIR, these models are often computationally expensive and complex. Their large size and high computational demands make them impractical for deployment on resource-constrained devices such as mobile phones or embedded systems. Striking a balance between performance and efficiency remains a significant issue, requiring models to maintain high restoration quality without becoming too cumbersome.
V-A4 Limited High-Quality Real-world Data
@@
Many AiOIR models depend on large-scale, high-quality labeled datasets for supervised training, but acquiring these datasets is resource-intensive. Real-world degraded data is often scarce, and the unpredictability of degradations makes it difficult for models to perform well in practical applications.
V-A5 High Dimensional Data
@@
Current IR models primarily focus on 2D images, but handling 3D data and video sequences presents additional challenges. For video restoration, not only does each frame need high-quality restoration, but temporal consistency across frames must also be preserved. This adds complexity and demands more sophisticated methods that can integrate spatial and temporal information simultaneously.
@@ V-B Future Research Directions
Future research will focus on several key directions. First, establishing a robust multi-task learning theory will be essential to effectively address task conflicts and optimize information-sharing mechanisms. Second, reducing reliance on large-scale labeled datasets by exploring semi-supervised and unsupervised learning methods will improve the model’s adaptability in data-scarce situations. Additionally, designing efficient models suitable for edge computing will make AiOIR more feasible for practical applications. Furthermore, investigating more complex real-world degradation scenarios will drive improvements in model performance across various environments. Finally, integrating large multimodal pretrained models and generative priors will enhance the restoration capabilities by leveraging rich multimodal data. These research directions will lay the foundation for the practicality and flexibility of AiOIR models, enabling them to tackle a broader range of real-world challenges.
V-B1 Incorporating Robust Multi-task Learning Theory
The development of multi-task learning (MTL) theory for AiOIR is still in its early stages, presenting key research opportunities [57, 263]. Challenges include managing task conflicts, dynamically weighting tasks, and optimizing inter-task information sharing to avoid interference [126, 127]. Further improvements can come from determining optimal task sequences, prioritizing tasks based on degradation severity, and designing loss functions that balance conflicting objectives. Progress in these areas is essential for building efficient, generalizable AiOIR systems for real-world degradations.
V-B2 Robust Prompt Learning under Data Scarcity
While prompt-based learning enhances adaptability in AiOIR, its tendency to overfit under limited or imbalanced training data is a key challenge. Future work may explore prompt-space regularization (e.g., sparsity or low-rank constraints) to limit over-specialization, as well as meta-learning frameworks (e.g., MAML) to promote generalization across unseen degradations. Incorporating degradation-aware conditioning can further improve robustness. Overall, designing prompt mechanisms that generalize reliably remains an open and important direction.
V-B3 Semi-Supervised and Unsupervised Learning Methods
@@
Reducing the dependency on large-scale labeled datasets is essential for the scalability and applicability of AiOIR models. Future research should focus on developing semi-supervised and unsupervised learning approaches that can learn effective representations from unlabeled or partially labeled data [264, 265]. Techniques such as self-supervised learning, contrastive learning, and unsupervised domain adaptation can be leveraged to improve model performance in scenarios where labeled data is scarce or unavailable. By advancing these methods, AiOIR models can become more adaptable to diverse and unforeseen degradations encountered in real-world conditions.
V-B4 Platform-Aware Model Design and Efficient Approaches
An important direction is developing edge-friendly AiOIR models that balance accuracy and efficiency for deployment on mobile and embedded devices [266, 267]. Techniques like compression, pruning, quantization, and architecture search help reduce resource use with minimal performance loss. In addition to these methods, more radical approaches such as binary neural networks (BNNs) and spiking neural networks (SNNs) offer promising avenues for AiOIR deployment. Recent works show their effectiveness in tasks like super-resolution and deraining [268, 269, 270, 271]. Combined with lightweight transformers, edge-specific distillation, and early-exit strategies, these methods advance practical AiOIR deployment.
V-B5 Addressing More Practical and Complex Degradations
@@
There is a need to focus on more practical tasks and datasets that reflect the complexities of real-world image degradations [272, 28]. Future research should explore composite and complex degradation scenarios, such as image restoration in nighttime conditions, deraining and dehazing in the dark, and images affected by multiple overlapping degradations rather than isolated, mixed single degradation tasks. Developing and utilizing datasets that capture these challenging conditions will enable models to learn from and be tested on data that closely resembles real-world challenges. This focus will drive the development of AiOIR models that are more robust and effective in practical applications.
V-B6 Incorporating Large Multimodal Pretrained Models and Generative Priors
@@
Another promising avenue is exploiting large multimodal pretrained models, particularly those incorporating generative models, to enhance AiOIR tasks [273, 274, 275, 276]. Models like CLIP [86] and recent advancements in generative models (e.g., stable diffusion [208]) have demonstrated remarkable abilities to capture complex data distributions across multiple modalities. By mining the rich representations and priors from these universal models, AiOIR can benefit from enhanced understanding of image content and context, leading to better restoration in low-level tasks.
V-B7 Leveraging Multi-Modal Information
@@
Most current AiOIR models primarily rely on single-modal image information, such as RGB images, limiting their effectiveness in addressing complex restoration tasks. Future research can focus on integrating multi-modal information—like depth maps, optical flow, and infrared images—into AiOIR frameworks [277, 278, 279]. This integration would provide models with rich contextual and structural insights, enhancing their ability to accurately recover images with varied degradations [280].
V-B8 Establishment of Standardized Evaluation Protocols and Benchmarks
Unlike single image restoration tasks with established datasets, AiOIR lacks standardized benchmarks. To enable fair comparison, it’s essential to develop unified evaluation protocols and diverse datasets covering real-world scenarios like high-resolution restoration, medical enhancement, old photo repair, and challenging weather conditions.
V-B9 Extension to High-dimensional Data
@@
In addition to RGB images, extending AiOIR methods to other data, such as video [281, 282], 3D data [282, 283], dynamic event data [284, 285], and hyperspectral data [286], presents significant opportunities for future research. Developing techniques that effectively integrate spatial, temporal, spectral, dynamic, and 3D information will be vital for applications such as video enhancement, 3D rendering, spectral analysis, and augmented reality. Addressing these challenges will expand the capabilities of AiOIR models, making them more versatile and applicable to a broader range of tasks.
V-B10 Privacy Protection
AiOIR is increasingly deployed in medical imaging, surveillance, and autonomous driving, yet its ethical implications are seldom addressed. Prior work on machine unlearning [135] shows how to erase the influence of private data in AiOIR, suggesting future research can explore data deletion and privacy-preserving training. Beyond privacy, biased data or model priors can cause fairness issues like uneven performance or hallucinations. Furthermore, misuse risks must be considered, as restored content may be used maliciously in misinformation, surveillance overreach, or unauthorized identity synthesis. These issues highlight the need for responsible development and deployment.
@@ VI Conclusion
In this paper, we present a comprehensive review of recent advancements in all-in-one image restoration, a rapidly emerging domain that integrates multiple types of image degradations into a single framework. Through an in-depth exploration of state-of-the-art models, we emphasize their robust capabilities, diverse architectures, and the rigorous experiments. By contrasting these models with traditional single-task approaches, we underscore the limitations of the latter in addressing real-world complexities, while highlighting the significant gains in efficiency, adaptability, and scalability offered by AiOIR models.
Our comprehensive taxonomy of existing works offers a multi-dimensional perspective, covering structural innovations, key methodologies such as prompt learning, mixture-of-experts, the incorporation of multimodal models, and autonomous agents. We further present an analysis of critical datasets, providing researchers and practitioners with a holistic toolkit to better assess the current landscape of AiOIR. Despite the considerable strides made in recent years, challenges persist. Current models still struggle with handling complex and compounded degradations, lack computational efficiency, and do not generalize well in real-world scenarios. We believe that future research will focus on several key areas: the development of more lightweight and efficient architectures, advancements in semi-supervised learning, and expanding the scope of AiOIR models to accommodate multimodal inputs and video data. Additionally, as the field progresses, innovations in cross-modal learning, real-time processing, and interpretability will likely become central to pushing the boundaries of what AiOIR systems can achieve.
In conclusion, AiOIR represents more than just a technical integration of multiple degradation types—it marks a paradigm shift toward unified, general-purpose vision models that can adapt to real-world complexity with minimal supervision. While significant progress has been made, substantial challenges remain, especially in scaling to real-world conditions, ensuring robustness under compounded degradations, and extending beyond visual-only inputs. We foresee future innovations at the intersection of efficient architecture design, multimodal representation learning, and autonomous adaptation mechanisms. By charting a comprehensive landscape of AiOIR, this review aims not only to inform but also to inspire a new wave of research that pushes the boundaries of what is possible in image restoration and beyond.